OpenStudy (anonymous):

Find the equation of the line whose slope is 0 and that contains the point (-4, 2). Answer x - 2 = 0 y + 4 = 0 y - 2 = 0 x + 4 = 0

OpenStudy (anonymous):

Stop posting question after question, especially identical ones, and try to understand what you're doing.

OpenStudy (ash2326):

bailey what's the slope of a horizontal line?

OpenStudy (anonymous):

either your going to help me or not, im going to continue to post and post b/c i don't understand Sooo...therefore if when i post the question and you want to elaborate then goo head ..... but @ ashh and I do not know ! explain it 2 me !

OpenStudy (anonymous):

is it y1-y2 over x1-x2 !?

OpenStudy (ash2326):

Bailey. Nae whenever anyone post just answer to your question , ask them to explain it. See we have the standard equation of a line in slope intercept form y=mx+c m= slope c= intercept on y-axis here slope m=0 so y=c the line contains the point (-4, 2) so 2=c hence y=c we found c=2 so y=2 or y-2=0

OpenStudy (anonymous):

Thanks (: ok soo on this one i do the same ? : Find the equation of the line that passes through (4, 5) with a slope of 2. Write your answer in Slope-Intercept Form. Answer y = 2x - 5 y = 2x + 1 y = 2x + 3 y = 2x - 3

OpenStudy (ash2326):

Yeah put m=2 and then satisfy the point (4, 5) in the equation to find c try it , if you get stuck I'll help you

OpenStudy (anonymous):

i got y=2x-3

OpenStudy (ash2326):

Correct, great job bailey:D

OpenStudy (anonymous):

Thank you (: && thanks soo muh for being patient !!!

OpenStudy (ash2326):

You're welcome bailey: D (: